OpenStudy (anonymous):

how do i solve this 5cosθ -2sinθ =2

OpenStudy (phi):

you can replace the sin (or cos) using sin^2 + cos^2 = 1

OpenStudy (xapproachesinfinity):

you can't divide since cos(theta) could be zero!

OpenStudy (xapproachesinfinity):

you can do \(25\cos^2\theta+4\sin^2\theta-20\sin\theta\cos\theta=4\) \(21\cos^2\theta-20\sin\theta\cos\theta=0\) take if from there

OpenStudy (xapproachesinfinity):

of course i used cos^2+sin^2=1 an simplified